OpenStudy (anonymous):

7/15 + 9/ 15 + 4/15 = ??

OpenStudy (anonymous):

20/15

OpenStudy (kendricklamar2014):

What have u done?

OpenStudy (anonymous):

4/3

OpenStudy (benlindquist):

common denominator right

OpenStudy (kendricklamar2014):

@AlexKross dont give direct answers. Its against the rules

OpenStudy (anonymous):

it there any way it can be a mixed number?

OpenStudy (benlindquist):

improper can be turned into mixed i think

OpenStudy (anonymous):

do you want me to give you the choices?

OpenStudy (kendricklamar2014):

9/15 can be simplified to 3/15 \[\frac{ 9 }{ 15 } = \frac{ 3 }{ 15 }\]

OpenStudy (benlindquist):

i thought people can't give answers

OpenStudy (kendricklamar2014):

Add 7/15 + 3/5:\[\frac{ 7 }{ 15 } + \frac{ 3 }{ 15 } = \frac{ 16 }{ 15 }\]

OpenStudy (benlindquist):

?? ✌🏻

OpenStudy (anonymous):

its multiple choice A) 7/9 B) 1 13/15 C) 2 2/9 D) 1 1/3

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (kendricklamar2014):

Can you draw the problem? @rainbowswirl

OpenStudy (anonymous):

what do you mean? @KendrickLamar2014

OpenStudy (kendricklamar2014):

Click on the button I circled in red: http://prntscr.com/9a1tr6

OpenStudy (anonymous):

now what do i do? @KendrickLamar2014

OpenStudy (kendricklamar2014):

You click the draw button and then draw the problem

OpenStudy (anonymous):

do you want me to draw the answers? @KendrickLamar2014

OpenStudy (kendricklamar2014):

Nope, I want you to draw the problem

OpenStudy (anonymous):

could you make up a problem to show me, I dont know how to add fractions :(

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (kendricklamar2014):



OpenStudy (kendricklamar2014):

@rainbowswirl

OpenStudy (anonymous):

I got 1/1?! @KendrickLamar2014

OpenStudy (anonymous):

would it be 1 13/15?

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (kendricklamar2014):

No you draw your problem like that ^